MATH 3023-Quiz # Name: Consider a shipment of computers. Historical data indicates that a shipment of this type has 5% of defects are with CPU only, 7% of the defects are with the RAM, and 3% of the defects are with both CUP and RAM. If a computer from this shipment is selected randomly, what is the probability that the computer has a) RAM only defects? b) neither CPU nor RAM defects?
Answer #1

a)P(RAM defects only)=P(RAM defects)-P(RAM and CPU both defects)=0.07-0.03=0.04

b)

P(neither CPU nor RAM)=1-P(at least one) =1-(P(RAM only)+P(CPU only)+P(RAM and CPU both)

=1-(0.04+0.05+0.03)=0.88